# Revealing Nanostructures in High-Entropy Alloys via Machine-Learning Accelerated Scalable Monte Carlo Simulation

### v1

@@ 4 Conclusions
    The GPU accelerated implementation of the SMC method enables the simulation of atomistic systems exceeding one billion atoms while maintaining the accuracy of density functional theory (DFT). Such unprecedented capability makes it possible for us to directly observe the nanostructures in HEAs, which can be of more than millions of atoms and are vital for the supurb mechanical properties in HEAs.
@@
    Using the SMC_GPU code, we investigated two types of high entropy alloys that have demonstrated exceptional mechanical properties. For the fcc FeCoNiAlTi alloy, we identified that at the experimental aging temperature, nanoparticles of sizes about 10-20 nm will form within the random matrix. The nanoparticles forms a slightly randomized $L1_{2}$ structure, with Ni and Co mainly occupies the A sites, and Al Ti mainly occupies the B sites. This size and composition of the nanoparticles agree excellently with the experimental findings [39]. Morover, we find that seemingly separate grains may, in fact, be connected, highlighting the intricate nature of high-entropy alloys (HEAs) and prompting a reconsideration of traditional grain size measurement methods.
@@
    We further investigated the bcc MoNbTaW high-entropy alloy (HEA) using a one-billion-atom supercell. Our results reveal the formation of hierarchical nanostructures. The smaller nanoparticles consist of a disordered A2 structure comprising Nb and W, with a feature size of approximately 10 nm, and are observed exclusively at low temperatures. In contrast, the larger nanophase exhibits a B2 structure composed of Ta and Mo, with a feature size of about 100 nm. Interestingly, we find that the Ta-Mo B2 structures decompose into two distinct nanophases, B2+ and B2-, which are separated by a grain boundary enriched with Nb and W. Simulations suggest that the disordered Nb-W grain boundary acts as a protective barrier for the two nanophases as the temperature increases, which is an interesting theoretical phenomenon to be confirmed in experiment in the future.
@@ 5.1 Ising model
Ising model is a fundamental model in statistical mechanics [72]. The study of Ising model not only helps elucidating a plethora of important physics phenomena, such as ferromagnetic phase transition, renormalization-group, and critical behaviour [5], but also leads to the invention of many widely-used algorithms, such as the Swendsen-Wang algorithm [73], Wang-Landau algorithm [74], and Wolff cluster flipping algorithm [75, 76]. In the simplest form, the Ising Hamiltonian can be written as:
@@
where $J$ is the interacting parameter and $\sigma_{i}$ represents the spin at the neighboring lattice site $i$, and $\bm{\sigma}$ denotes a spin configuration in the lattice. Note that the summation is over all sites and their nearest neighbors. For example, in 2D square lattice, there are four nearest neighbors, and in 3D square lattice there are 6 nearest neighbors. It is also straightforward to generalize the above model to include longer-distance pair interactions within a predefined cutoff radius $r_{c}$ in a general lattice. Such Ising-like models can then be applied to study real materials, in which the spin variable $\sigma$ represents the different elements, and the generalized interaction parameters $J_{\sigma_{i}\sigma_{j}}$ can be interpreted as the chemical bonding between atoms of specific elements.
@@ 5.3 Monte Carlo Simulation Algorithms
A sample following the probability distribution of Eq. 8 can be generated via Markov chain Monte Carlo (MCMC) simulation. For such a purpose, different schemes can be devised, and the Metropolis-Hastings algorithms is among the most commonly used ones. For the Ising model in Eq. 7, the Metropolis-Hasting algorithm starts with proposing flipping the spins of each lattice site. This trial move gives rise to a change of the total energy by $\Delta E$, and the acceptance probability of this trial is given by:
@@
The Monte Carlo simulation can be carried out at different conditions, the most common ones are the canonical ensembles, grand-canonical ensembles, and semi-grand canonical ensembles. In canonical ensemble, the number of particles and the temperature is kept fixed. In the grand-canonical ensemble, the chemical potential is kept fixed, and the number of particles is determined by the equilibrium condition. The semi-grand canonical is a technique to fix the concentration of different elements by adjusting the value of the chemical potentials of each element. In this work, we focus on the canonical ensemble, which is relatively straightforward to implement. In a canonical Monte Carlo simulation, the first thing to note is that the flipping trial in Ising model is replaced with a swapping trial, in which a random pair of sites are chosen randomly to switch the particles.
@@ 5.4 Parallelization opportunities in Monte Carlo simulation
    Temperature parallelism: the simulations at different temperatures are independent, therefore they are “embarrassingly parallel”. A closely related case is the replica exchange method, in which only the temperature of different replica need to be exchanged.
@@
    Sub-move parallelism: For calculating the energy change due to a single MC move, some fine-grain parallelization opportunities usually exist, depending on the form of the effective Hamiltonian. This is particularly true for the models based on neural networks, where the large number of matrix-multiplication operations present excellent opportunities for acceleration with highly optimized math libraries.
In this work, we focus our discussion on move-parallelism, since it is the bottleneck for large-scale MC simulation. Implementation of the temperature parallelism is straightforward. For sub-move parallelism, we make use of the fact that, in the general case, all the local energies of sites within a LIZ need to be evaluated, and the calculations of these local energies are independent of each other.

### v2

@@ 4 Conclusions
    The GPU-accelerated implementation of the SMC method enables the simulation of atomistic systems exceeding one billion atoms while maintaining the accuracy of density functional theory (DFT). Such unprecedented capability makes it possible for us to directly observe the nanostructures in HEAs, which can be of more than millions of atoms. Understanding the nanoscale evolution of NPs are vital for understanding the origin of the superb mechanical properties in HEAs.
@@
    Using the SMC-GPU code, we investigated the size, composition, and morphologies of the NPs in the fcc $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA using a one-billion-atom supercell. This large supercell not only makes it possible to directly observe the nanostructure evolution in HEAs, but also enables the direct comparison with experimental results from TEM and APT. This size and composition of the nanoparticles generally align well with available experimental findings [39]. Moreover, we find that seemingly separate NPs may, in fact, be connected, highlighting the intricate nature of high-entropy alloys (HEAs) and prompting a reconsideration of traditional grain size measurement methods.
@@
    We further investigated the bcc MoNbTaW high-entropy alloy (HEA) using a one-billion-atom supercell. The results reveal the formation of hierarchical nanostructures: The smaller NPs consist of a disordered A2 structure comprising Nb and W, with a feature size of approximately 10 nm; The larger nanophases exhibits a B2 structure composed of Ta and Mo, with a feature size of about 100 nm. Interestingly, we find that the Ta-Mo B2 structures decompose into two distinct nanophases, B2+ and B2-, which are separated by a grain boundary enriched with Nb and W. The simulation results suggest that the disordered Nb-W grain boundary acts as a protective barrier for the two nanophases to reduce the enthalpy of the system.
@@ 5.1 Ising model
Ising model is a fundamental model in statistical mechanics [75]. The study of the Ising model not only helps elucidating a plethora of important physics phenomena, such as ferromagnetic phase transition, renormalization-group, and critical behaviour [5], but also leads to the invention of many widely-used algorithms, such as the Swendsen-Wang algorithm [76], Wang-Landau algorithm [77], and Wolff cluster flipping algorithm [78, 79]. In the simplest form, the Ising Hamiltonian can be written as:
@@
where $J$ is the interacting parameter and $\sigma_{i}$ represents the spin at the neighboring lattice site $i$, and $\bm{\sigma}$ denotes a spin configuration in the lattice. Note that the summation is over all sites and their nearest neighbors. For example, in a 2D square lattice, there are four nearest neighbors, and in a 3D square lattice, there are 6 nearest neighbors. It is also straightforward to generalize the above model to include longer-distance pair interactions within a predefined cutoff radius $r_{c}$ in a general lattice. Such Ising-like models can then be applied to study real materials, in which the spin variable $\sigma$ represents the different elements, and the generalized interaction parameters $J_{\sigma_{i}\sigma_{j}}$ can be interpreted as the chemical bonding between atoms of specific elements.
@@ 5.3 Monte Carlo Simulation Algorithms
A sample following the probability distribution of Eq. 8 can be generated via Markov chain Monte Carlo (MCMC) simulation. For such a purpose, different schemes can be devised, and the Metropolis-Hastings algorithm is among the most commonly used ones. For the Ising model in Eq. 7, the Metropolis-Hasting algorithm starts with proposing flipping the spins of each lattice site. This trial move gives rise to a change of the total energy by $\Delta E$, and the acceptance probability of this trial is given by:
@@
The Monte Carlo simulation can be carried out at different conditions, the most common ones are the canonical ensembles, grand-canonical ensembles, and semi-grand canonical ensembles. In a canonical ensemble, the number of particles and the temperature are kept fixed. In the grand-canonical ensemble, the chemical potential is kept fixed, and the number of particles is determined by the equilibrium condition. The semi-grand canonical is a technique to fix the concentration of different elements by adjusting the value of the chemical potentials of each element. In this work, we focus on the canonical ensemble, which is relatively straightforward to implement. In a canonical Monte Carlo simulation, the first thing to note is that the flipping trial in the Ising model is replaced with a swapping trial, in which a random pair of sites is chosen randomly to switch the particles.
@@ 5.4 Parallelization opportunities in Monte Carlo simulation
    Temperature parallelism: the simulations at different temperatures are independent, therefore they are “embarrassingly parallel”. A closely related case is the replica exchange method, in which the configurations from different temperatures are constantly exchanged in the simulation process, with a policy satisfying detailed balance condition, in order to enhance the speed for reaching thermal equilibrium.
@@
    Sub-move parallelism: For calculating the energy change due to a single MC move, some fine-grain parallelization opportunities exist, depending on the form of the effective Hamiltonian. This is particularly true for the models based on neural networks, where a large number of matrix-multiplication operations present excellent opportunities for acceleration with highly optimized math libraries.
In this work, we mainly focus on move-parallelism, which is one of the bottleneck for large-scale MC simulations. For sub-move parallelism, we make use of the fact that, in the general case, all the local energies of sites within a LIZ need to be evaluated, and the calculations of these local energies are independent of each other.